How many liters of a 25% antifreeze solution must be mixed with 4 liters of a 60% antifreeze solution to make a mixture that is 45% antifreeze?
Answer by richwmiller(17219)

0.25*x+0.6*4=0.45(4+x)
0.25x-0.45x=1.8-2.4
-0.2x=-0.6
x=3